# Towards Interpretable Sequence Continuation: Analyzing Shared Circuits in Large Language Models

### v5

@@ Abstract
While transformer models exhibit strong capabilities on linguistic tasks, their complex architectures make them difficult to interpret. Recent work has aimed to reverse engineer transformer models into human-readable representations called circuits that implement algorithmic functions. We extend this research by analyzing and comparing circuits for similar sequence continuation tasks, which include increasing sequences of Arabic numerals, number words, and months. By applying circuit interpretability analysis, we identify a key sub-circuit in both GPT-2 Small and Llama-2-7B responsible for detecting sequence members and for predicting the next member in a sequence. Our analysis reveals that semantically related sequences rely on shared circuit subgraphs with analogous roles. Additionally, we show that this sub-circuit has effects on various math-related prompts, such as on intervaled circuits, Spanish number word and months continuation, and natural language word problems. Overall, documenting shared computational structures enables better model behavior predictions, identification of errors, and safer editing procedures. This mechanistic understanding of transformers is a critical step towards building more robust, aligned, and interpretable language models.
@@ 1 Introduction
The main contributions of this work are:
@@
These contributions improve the understanding of how transformer models generalize concepts by re-using components for analogous tasks.
Figure 1: The important components of a shared, entangled sub-circuit for the Numerals, Number Words, and Months tasks in GPT-2 Small. The functional roles of the components are labeled below them. Resid_post denotes the residual stream state right before the linear unembedding to logits. Full circuits are shown in Appendix H.
@@ 2 Background and Related Work
Another type of causal intervention is activation patching, which takes the corrupted dataset as input, and then restores the activations at a certain component with the original activations to observe how much that restored component recovers the original performance. Path patching is a different type of patching that allows for a more precise analysis of an intervention’s effect on a particular path Goldowsky-Dill et al. (2023). It can be performed by ablating component interactions, measuring the effect of one component on another. The Automatic Circuit DisCovery (ACDC) technique employs iterative patching automatically find circuit graphs for given tasks (Conmy et al., 2023); however, this technique only seeks to automate finding the connectivity of circuit graphs, and not their functionality interpretation.
@@ 3.1 Connectivity Discovery Methods
Metrics. We utilize the logit difference to measure model task capability by taking the difference between the correct token $L_{C}$ and an incorrect token logit $L_{I}$. The incorrect logit may be chosen as a token that is not the correct token. To compare an ablated model with the unablated model, we employ the performance score, a percentage calculated as the logit difference of the ablated model over the logit difference of the unablated model.
Iterative Pruning for Nodes. To search for circuit components, we use a knockout method that ablates one candidate component (node) at a time and checks how much performance falls. This method begins with all the components as a candidate circuit. At each step, ablation is performed by patching in the mean activations of a corrupted dataset at a candidate node, plus all the nodes not in our candidate circuit. If performance falls below $T_{n}$, a user-defined performance threshold, the node is kept for the candidate circuit, as it is deemed necessary for the task. Else, it is removed.
We start by removing components from the last layer, continuing until the first layer; we call this procedure the backward sweep. At each layer during the backward sweep, we first ablate the layer’s MLP, and then consider its attention heads. Next, we then prune again from the first layer to the last layer; we call this the forward sweep. At each layer during the forward sweep, we first ablate each attention heads, and then its MLP. We continue iterating by successive backward-forward sweeps, stopping when no new components are pruned during a sweep. The output is the unpruned node set.
@@
Iterative Path Patching for Edges. After finding circuit nodes, we utilize path patching to obtain interactions (edges) between them. Edges denote nodes with high effects on other nodes  [^3] As the residual stream allows for indirect effects, edges may be between components at non-adjacent layers,. We apply a form of iterative path patching which works backwards from the last layers by finding earlier components that affect them. First, we ablate the nodes pruned from iterative node pruning. Then, we ablate one candidate edge of the unablated nodes at a time. Using the same order as the backward sweep, we take a node as a receiver and find the sender nodes that have an important effect on it. If patching the effect of sender A on receiver B causes the model performance to fall below threshold $T_{e}$, the edge is kept; else, it is removed.
@@ 3.2 Functionality Discovery Methods
Attention Pattern Analysis. We analyze the $\mathbf{QK}$ matrix of attention heads to track information movement from keys to queries. We take the mean of datasets samples to calculate the attention scores of the heatmaps in §5.2 and §L, and display only one sample on the axes for demonstration.
Component Output Scores. We analyze head outputs by examining the values written to the residual stream via the heads’ output matrices ($\mathbf{OV}$), allowing us to see what information is being passed by each head along in the circuit. These values are measured by component output scores; we utilize a "next sequence" score that measures how well a head, given sequence token $I$, outputs token $I+1$. The details of this method are described in Appendix K.
@@ 4.1 Experimental Setup
In this section, we describe the experimental setup for our ablation experiments. We ran experiments on a NVIDIA A100 GPU. More setup details can be found in Appendix F.
Models. We test on GPT-2 Small (117M parameters), which has 144 heads and 12 MLPs
@@
Datasets. For GPT-2 Small, we run a generated prompts dataset of length 4 sequences (eg. 1 2 3 4). Our focus in this paper is not on finding circuits only for Numeral sequences, but on prompt types that share a common abstraction. Thus, to better compare numbers to months, we use sequences ranging from 1 to 12.
For each task, we generate samples by placing our sequence members among non-sequence tokens. For instance, one sample may be ’Kyle was born in February… Grant was born in April. Madison was born in’. Placing sequence members amongst non-sequence tokens allows us to evaluate the circuit representation of the shared sub-task of how the model selects sequence members from non-sequence members. We generate a total of 1536 samples per task, with a total of 4608 samples.
@@
Metric. We measure using logit difference using the last sequence member as the incorrect token (eg. 1 2 3 has 4 as correct, and 3 as incorrect).
@@ 4.2 Shared Sub-Circuits for Similar Sequence Continuation Tasks
In this section we analyze main results, and include extended results in Appendix §B. We discover shared sub-circuits across the three sequence continuation tasks. These circuits were found using a performance threshold of $T_{n}=T_{e}=80$%. Figure 1 shows a sub-circuit found across the circuits for all three tasks, which includes heads 4.4, 7.11 and 9.1, which we show to be important in Table 1. Figure 9 in Appendix H combines all three circuits into one graph  [^5] Due to the (qkv) circuit’s large display size, we show the circuits with (qkv) decomposition in Appendix I..
Several important attention heads are identified across various circuits. We define a head as important if their ablation from a circuit causes an average drop of at least -20% performance for all tasks 66620% is chosen due to using $T_{n}=80\%$, so that for many removal order variations, a component with a 20% importance cannot be removed, unless there are alternative backups.. Table 1 compares the importance of these attention heads for our tasks. We note that ablating heads 0.1, 4.4, 7.11, and 9.1 cause drops >20% for all three circuits, while ablating 1.5 causes a drop >20% for Numerals and Number Words circuits.
Table 1: Drop in Task Performance when a Head is Removed from a Circuit in Figure 9.
@@
MLP Connectivity. For all tasks, we find that several MLP ablations cause a >20% performance drop. In particular, MLP 9 causes a substantial drop of more than 90%. These results are found in Appendix J.
@@ 5 Explaining Shared Component Functionalities in GPT-2 Small
In this section we analyze main results, and include extended results in Appendix §C.
@@ 5.1 Sub-Circuit Hypothesis
We hypothesize how the important shared components for the three tasks work together as a functional sub-circuit. We define sub-tasks that all three sequence continuation tasks share: (1) Identifying Sequence Members and (2) Predicting the Next Member after the Most Recent Member.
@@
Our hypothesis is that early heads, in particular 1.5 and 4.4, identify similar, adjacent sequence members, such as numbers or Months, without yet attending to the distinction of which numbers should be focused on more than others. Following this, information is passed further along the model to heads, such as 7.11, to discern consecutive number sequences and deem the two most recent elements as more significant. This information is then conveyed to head 9.1 to put more emphasis on predicting the next element in the sequence. Lastly, the next element calculation is done primarily by MLP 9. Thus, this sub-circuit would represent an algorithm that carries out the sub-tasks shared for all three tasks. This section details evidence that supports this circuit hypothesis.
5.2 Attention Head Functionality
5.2.1 Sequence Member Detection Heads
@@
Figure 2: GPT-2 Small attention patterns for (a) Head 1.5 and (b) Head 4.4. Lighter colors mean higher attention values. For each of these detection patterns, the query is shown in green, and the key is shown in blue. The Months are in sequential order, but the Numerals are not. For head 1.5, similar types attend to similar types. But for head 4.4, Months attend to Months, but Numerals do not attend to Numerals. For all plots, we take the mean of dataset samples to calculate the attention scores, but display only one sample on the axes for demonstration purposes.
@@
We discover a "similar member" detection head 1.5, and a "sequence member detection" 4.4. Attention pattern analysis reveals that these heads detect how sequence members (as queries) attend to sequence members (as keys) of the same type, such as numerals. To determine if this detection only occurs if the sequence members are in sequential order, or if this occurs even if they are not, we input prompts with Numerals in random order but with Months in sequential order. In Figure 2, we observe, for head 1.5, similar types attend to similar types. However, for head 4.4, Months attend to Months, but Numerals do not attend to Numerals, as the Numerals are not in sequential order. Therefore, in general, both heads 1.5 and 4.4 appear to detect similar token types that belong to an ordinal sequence such Numerals or Months, but head 4.4 acts even more specifically as an adjacent sequence member detection head. More discussion about these attention patterns are in Appendix L.
@@
Last Token Sequence Detection Head. In Appendix §C, we describe results that show how 7.11 acts to transfer the detected sequence information to the last token, which aids head 9.1 and MLP 9 in determining the successor member.
@@
Figure 3: This GPT-2 Small attention pattern for Head 9.1 shows the component pays strong attention to only the most recent sequence member.
5.2.2 Successor Components
@@
Figure 1 shows that head 9.1 receives information from both 4.4 and 7.11. 9.1, shown in Figure 3, pays strong attention to only the last member of the sequence, and it appears to attend even stronger to the last member than 7.11, as shown in Figure 5. OV score analysis determines that 9.1 is a successor head Gould et al. (2023).  [^7] Successor heads were found concurrently via independent discovery by Gould et al. (2023).
@@
Succession via MLP For most samples of all sequence types, logit lens reveals that the model does not predict the correct answer before MLP 9. However, after the information is processed through MLP 9, the model outputs the next sequence member. These findings suggests that MLP 9 is largely responsible for finding the next sequence member, even more so than Head 9.1, which may just be boosting this information and/or acting as a backup component. Logit lens results are in Appendix §J.
@@ 6 Interpreting Sequence Continutation in Llama-2-7B
We apply the same iterative methods from GPT-2 to discover important components for each task in Llama-2, which has 1024 heads and 32 MLPs, and find that these have similar functionalities to GPT-2 components. We discover these components to be: 5.25, 16.0, and 20.17.
Given that we can find analogues of this sub-circuit’s components from one model to another, we show that this sub-circuit type can appear in more than one model. Future studies can examine if this appears in multiple models.
Shared Components: Attention Patterns. As shown in Figure 4, Llama-2-7B’s 5.25 appears to be a sequence detection head similar to GPT-2 Small’s 4.4. Moreover, it is not just a similar member detection head like GPT-2 Small’s 1.5; when sequence members are out-of-order, the head does not detect these members. But when the members of any of the 3 tasks are in-order, it is able to detect them. Likewise, in Figure 6, Llama-2-7B’s 20.17 appears to be a successor head similar to GPT-2 Small’s 9.1. We discuss 16.0 in Appendix 6.
Figure 4: The attention pattern of Head 5.25 for Llama-2-7b resembles the attention pattern of GPT-2 Small’s Head 4.4 in Figure 2(b), indicating they have similar functionality as sequence member detection heads.
Shared Components: OV Scores. 20.17 has a higher than average successor head score, as it is one of only twelve attention heads (out of 1024 total) which detects the "next" token of a sequence member token.
Spanish Tasks Functionality. For Spanish number words and months, we obtain similar results for attention patterns and OV scores as their English task counterparts.

### v6

@@ Abstract
While transformer models exhibit strong capabilities on linguistic tasks, their complex architectures make them difficult to interpret. Recent work has aimed to reverse engineer transformer models into human-readable representations called circuits that implement algorithmic functions. We extend this research by analyzing and comparing circuits for similar sequence continuation tasks, which include increasing sequences of Arabic numerals, number words, and months. By applying circuit interpretability analysis, we identify a key sub-circuit in both GPT-2 Small and Llama-2-7B responsible for detecting sequence members and for predicting the next member in a sequence. Our analysis reveals that semantically related sequences rely on shared circuit subgraphs with analogous roles. Additionally, we show that this sub-circuit has effects on various math-related prompts, such as on intervaled circuits, Spanish number word and months continuation, and natural language word problems. This mechanistic understanding of transformers is a critical step towards building more robust, aligned, and interpretable language models.
@@ 1 Introduction
The main contributions of this work are [^1] To encourage reuse and further development our code and datasets can be found here: https://github.com/apartresearch/seqcont_circuits:
@@
Figure 1: The important components of a shared, entangled sub-circuit for the Numerals, Number Words, and Months tasks in GPT-2 Small. The functional roles of the components are labeled below them. Resid_post denotes the residual stream state right before the linear unembedding to logits. Full circuits are shown in Appendix J.
@@ 2 Background and Related Work
Another type of causal intervention is activation patching, which takes the corrupted dataset as input, and then restores the activations at a certain component with the original activations to observe how much that restored component recovers the original performance. The Automatic Circuit DisCovery (ACDC) technique employs iterative patching automatically find circuit graphs for given tasks (Conmy et al., 2023); however, this technique only seeks to automate finding the connectivity of circuit graphs, and not their functionality interpretation.
@@ 3.1 Connectivity Discovery Methods
Metrics. We utilize the logit difference to measure model task capability by taking the difference between the correct token $L_{C}$ and an incorrect token logit $L_{I}$. The incorrect logit may be chosen as a token that is not the correct token. To compare an ablated model with the unablated model, we employ the performance score, a percentage calculated as the logit difference of the ablated model over the logit difference of the unablated model. We define an important component to be a component that, when ablated, greatly reduces performance.
Iterative Pruning for Nodes. To search for circuit components, we use a knockout method that ablates one candidate component, or node, $n_{c}$ at a time and checks how much performance falls. This method begins with all the components as a set $C$, which are nodes of a candidate circuit. Let $N$ denote the set of all the nodes in the network. At each step, ablation is performed by patching in the mean activations of a corrupted dataset at $N\setminus C$, which are the nodes that are not part of of the current candidate circuit. Then, the mean activations of the corrupted dataset are also patched in at new candidate node $n_{c}$. If performance falls below $T_{n}$, a user-defined performance threshold, node $n_{c}$ is kept in the candidate circuit $C$, as it is deemed necessary for the task. Else, it is removed.
We start by selecting component $n_{c}$ from the last layer, continuing until the first layer; we call this procedure the backward sweep. At each layer during the backward sweep, we first ablate the layer’s MLP, and then consider its attention heads. Next, we then prune again from the first layer to the last layer; we call this the forward sweep. At each layer during the forward sweep, we first ablate each attention heads, and then its MLP. We continue iterating by successive backward-forward sweeps, stopping when no new components are pruned during a sweep. The output is $C$, the unpruned node set.
@@
Iterative Path Patching for Edges. We describe this part of the method in Appendix D.
@@ 3.2 Functionality Discovery Methods
Attention Pattern Analysis. We analyze the $\mathbf{QK}$ matrix of attention heads to track information movement from keys to queries. We take the mean of datasets samples to calculate the attention scores of the heatmaps in this paper, and display only one sample on the axes for demonstration purposes.
Component Output Scores. We analyze head outputs by examining the values written to the residual stream via the heads’ output matrices ($\mathbf{OV}$), allowing us to see what information is being passed by each head along in the circuit. These values are measured by component output scores; we utilize a successor score that measures how well a head, given sequence token $I$, outputs token $I+1$. The details of this method are described in Appendix E.
@@ 4.1 Experimental Setup
We test on GPT-2 Small, which has 144 heads and 12 MLPs, with a total of 117M parameters.
@@
Datasets. We run a generated prompts dataset of length 4 sequences (eg. 1 2 3 4). Our focus in this paper is not on finding circuits only for Numeral sequences, but on prompt types that share a common abstraction. Thus, to better compare numbers to months, we use sequences ranging from 1 to 12.
For each task on GPT-2 Small, we generate samples by placing our sequence members among non-sequence tokens. For instance, one sample may be ’Kyle was born in February. Bob was born in March. Grant was born in April. Madison was born in’. Placing sequence members amongst non-sequence tokens allows us to evaluate the circuit representation of the shared sub-task of how the model selects sequence members from non-sequence members. We generate a total of 1536 samples per task, with a total of 4608 samples. More details can be found in Appendix B.
@@
Metric. We measure using logit difference using the last sequence member as the incorrect token (eg. for "1 2 3", the correct token is "4", the incorrect token is "3").
@@ 4.2 Shared Sub-Circuits for Similar Sequence Continuation Tasks
In this section we analyze main results, and include extended results in Appendix §F. We discover shared sub-circuits across the three sequence continuation tasks. These circuits were found using a performance threshold of $T_{n}=T_{e}=80$%. Figure 1 shows a sub-circuit found across the circuits for all three tasks, which includes attention heads 4.4, 7.11 and 9.1, which we show to be important in Table 1. Figure 8 in Appendix J combines all three circuits into one graph  [^4] Due to the (qkv) circuit’s large display size, we show the circuits with (qkv) decomposition in Appendix K..
Several important attention heads are identified across various circuits. We define an attention head as important if their ablation from a task’s circuit causes an average performance drop of at least 20% for all tasks 55520% is chosen due to using $T_{n}=80\%$, so that for many removal order variations, a component with a 20% importance cannot be removed, unless there are alternative backups.. Table 1 compares the importance of these attention heads for our tasks. We note that ablating attention heads 0.1, 4.4, 7.11, and 9.1 cause drop of more than 20% for all three circuits, while ablating attention head 1.5 causes a drop of approximately 20% for Numerals and Number Words circuits.
Table 1: Drop in Task Performance for GPT-2 Small when an important attention head is removed from the circuit in Figure 8.
@@
MLP Connectivity. For all tasks, we find that several MLP ablations cause a >20% performance drop. In particular, MLP 9 causes a substantial drop of more than 90%. These results are found in Appendix L.
@@ 5 Explaining Shared Component Functionalities in GPT-2 Small
In this section we analyze main results, and include extended results in Appendix §G.
@@ 5.1 Sub-Circuit Hypothesis
We hypothesize the important shared components for the three tasks work together as a functional sub-circuit. We define sub-tasks that all three tasks share: (1) Identifying Sequence Members and (2) Predicting the Next Member.
@@
Our hypothesis is that early attention heads, in particular 1.5 and 4.4, identify similar, adjacent sequence members, such as numbers or Months, without yet attending to the distinction of which numbers should be focused on more than others. Following this, information is passed further along the model to other components such as attention head 7.11, which may discern that the two most recent elements are more significant. This information is then conveyed to attention head 9.1 to put more emphasis on predicting the next element in the sequence. Lastly, the next element calculation is done primarily by MLP 9. Thus, this sub-circuit would represent an algorithm that carries out the sub-tasks shared for all three tasks. This section details evidence that supports this circuit hypothesis.
5.2 Sequence Member Detection Heads
@@
Figure 2: GPT-2 Small attention patterns for (a) Attention Head 1.5 and (b) Head 4.4. Lighter colors mean higher attention values. For each of these detection patterns, the query is shown in green, and the key is shown in blue. The Months are in sequential order, but the Numerals are not. For attention head 1.5, similar types attend to similar types. But for head 4.4, Months attend to Months, but Numerals do not attend to Numerals. For all plots, we take the mean of dataset samples to calculate the attention scores, but display only one sample on the axes for demonstration purposes.
@@
We discover a "similar member" detection attention head 1.5, and a "sequence member detection" attention head 4.4. Attention pattern analysis reveals that these heads detect how sequence members (as queries) attend to sequence members (as keys) of the same type, such as numerals. To determine if this detection only occurs if the sequence members are in sequential order, or if this occurs even if they are not, we input prompts with Numerals in random order but with Months in sequential order.
In Figure 2, we observe, for attention head 1.5, similar types attend to similar types. However, for attention head 4.4, Months attend to Months, but Numerals do not attend to Numerals, as the Numerals are not in sequential order. Therefore, in general, both attention heads 1.5 and 4.4 appear to detect similar token types that belong to an ordinal sequence such Numerals or Months, but attention head 4.4 acts even more specifically as an adjacent sequence member detection head. More discussion about these attention patterns are in Appendix M.
@@
Last Token Sequence Detection Head. In Figure 3 and in Appendix §G, we describe results that show how attention head 7.11 acts to transfer the detected sequence information to the last token, which aids attention head 9.1 and MLP 9 in determining the successor member.
@@
Figure 4: This GPT-2 Small attention pattern for Attention Head 9.1 shows that for the last token, the component pays strong attention to only the most recent sequence member.
5.3 Successor Components
@@
Figure 1 shows that attention head 9.1 receives information from both attention heads 4.4 and 7.11. Attention head 9.1, shown in Figure 4, pays strong attention to only the last member of the sequence, and it appears to attend even stronger to the last member than attention head 7.11, as shown in Figure 3. Attention head 9.1 has a successor score of 87.37 for a dataset of numerals from 1 to 97; more details on these scores is given in Appendix G.2. These results suggest that attention head 9.1 is a successor head Gould et al. (2023).  [^6] Successor heads were found concurrently via independent discovery by Gould et al. (2023).
@@
Succession via MLP For most samples of all sequence types, logit lens reveals that the model does not predict the correct answer before MLP 9. However, after the information is processed through MLP 9, the model outputs the next sequence member. These findings suggests that MLP 9 is largely responsible for finding the next sequence member, more so than attention head 9.1, which may just be boosting information and/or acting as a backup component. Logit lens results are in Appendix §L.
@@ 6 Interpreting Sequence Continutation in Llama-2-7B
We apply iterative pruning using datasets of sequences (without non-sequence words in between) to Llama-2-7B, which has 1024 attention heads and 32 MLPs, to find important attention heads 5.25, 16.0, and 20.17  [^7] For Llama-2-7B, we define a component as important is their ablation from a task’s circuit causes an average performance drop of at least 12% for all tasks. Specifically, attention head 5.25 is found to be a sequence detection head like GPT-2 Small’s attention head 4.4, and attention head 20.17 is found to be a successor head like GPT-2 Small’s attention head 9.1. Given that we can find analogues of the sub-circuit’s components described in §5.1 from one model to another, we show that this sub-circuit type appears in more than one model  [^8] Future studies can examine if this appears in even more models.. Appendix C details the setup of these experiments.
Important Components. As the full circuits’ component sets are between 80 to 100 nodes and are too large to include in this paper, we only discuss important components. Unlike in GPT-2 Small, for Llama-2-7B we find most of the MLPs to be necessary to be able to perform sequence continuation. Table 2 shows the drop in task performance for Llama-2-7B when an important attention head is removed from each respective task’s circuit.
Table 2: Drop in task performance for Llama-2-7B when an important attention head is removed from each sequence continuation task’s circuit.
Important Head Numerals NumWords Months
5.25 -29.24% -21.32% -12.87%
16.0 -27.02% -14.18% -13.74%
20.17 -56.78% -40.03% -27.26%
Attention Patterns. As shown in Figure 5, Llama-2-7B’s attention head 5.25 appears to be a sequence detection head similar to GPT-2 Small’s attention head 4.4. Moreover, it is not just a similar member detection head like GPT-2 Small’s attention head 1.5; when sequence members are out-of-order, the component does not detect these members. But when the members of any of the 3 tasks are in-order, it is able to detect them. Likewise, in Figure 6, Llama-2-7B’s attention head 20.17 appears to be a successor head similar to GPT-2 Small’s attention head 9.1. We discuss attention head 16.0 in Appendix 6.
Figure 5: The attention pattern of attention head 5.25 in Llama-2-7b resembles the attention pattern of GPT-2 Small’s attention head 4.4 in Figure 2(b), indicating they have similar functionality as sequence member detection heads.
OV Scores. Attention head 20.17 has a higher than average successor head score of 33.3%, compared to the average score of 0.46%. It is one of only twelve attention heads (out of 1024) which detects the "next" token of a sequence.
Figure 6: Llama-2-7B attention pattern for Head 20.17, bearing resemblence to the attention pattern of GPT-2 Small’s Head 9.1 in Figure 4, showing that for the last token, the component pays strong attention to only the most recent sequence member.